A sample containing NaBr and KBr only weighs 255 mg.  The sample was dissolved in water and treated with excess AgNO3.  The precipitate formed was found to weigh 420 mg.  Calculate the % NaBr in the sample. 


1
Expert's answer
2021-04-28T02:46:07-0400

The molar mass of NaBr is 102.894 g/mol.

The molar mass of KBr is 119.002 g/mol.

The molar mass of AgBr is 187.77 g/mol.


Solution:

If the solution of the mixture of NaBr and KBr is treated with AgNO3, an AgBr precipitate will be formed.

The chemical reactions for the formation of AgBr are written as follows:

NaBr + AgNO3 → AgBr + NaNO3

KBr + AgNO3 → AgBr + KNO3


Let, mass of NaBr = a, mass of KBr = b.

Here, mass of NaBr + KBr = 255 mg

Thus, a + b = 255


The amount of AgBr formed from NaBr is:

Mass of AgBr from NaBr = Mass of NaBr × (Molar mass of AgBr / Molar mass of NaBr) × (1/1)

Mass of AgBr from NaBr = (a) × (187.77 g mol-1 / 102.894 g mol-1) × (1/1) = 1.8249a

Mass of AgBr from NaBr = 1.8249a

The amount of AgBr formed from KBr is:

Mass of AgBr from KBr = Mass of KBr × (Molar mass of AgBr / Molar mass of KBr) × (1/1)

Mass of AgBr from KBr = (b) × (187.77 g mol-1 / 119.002 g mol-1) × (1/1) = 1.5779b

Mass of AgBr from KBr = 1.5779b


Here, mass of AgBr = 420 mg

Thus, 1.8249a + 1.5779b = 420


Let's consider the two equations:

a + b = 255

1.8249a + 1.5779b = 420


Multiply the first equation by -1.5779 and then add it to the second equation, we have

-1.5779a - 1.5779b = -402.3645

1.8249a + 1.5779b = 420

¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯¯

0.247a = 17.6355

a = 71.3988 = 71.4

Thus, mass of NaBr = a = 71.4 mg


%NaBr = (Mass of NaBr / Mass of sample) × 100%

%NaBr = (71.4 mg / 255 mg) × 100% = 28.0%

%NaBr = 28.0%


Answer: %NaBr in the sample is 28.0%.